# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

### v1

@@ 7 Case Studies and Examples
The use of AI in agriculture has enormous potential to revolutionize the way farming is conducted, from improving crop management to reducing costs and simplifying tasks for farmers. This potential is being realized by a number of start-ups and applications that leverage AI to address these challenges and improve productivity and efficiency in the agricultural sector. Here are some examples of how AI is being practically utilized in agriculture:
@@
  • Kisan GPT: It, launched on March 15, 2023, is an AI chatbot that caters to the underserved agriculture sector of India. It provides guidance on various topics such as irrigation, pest control, and crop cultivation. Unlike other popular AI chatbots where users are required to type their queries, Kisan GPT offers the option of voice input using a microphone. It supports 10 languages including Hindi, English, Gujarati, and Marathi, as well as foreign languages like Japanese, Portuguese, and Indonesian. With India being an agricultural powerhouse, this AI chatbot could be a game-changer for the agriculture sector. Farmer-friendly and accessible, it bridges the information gap between farmers and experts, aiding decision-making and enhancing crop management. With AI-powered chatbots, like Kisan GPT, farmers can achieve greater efficiency and make informed choices, benefiting India’s agricultural prowess [125].
@@
  • Jiva: Jiva provides a range of services, including micro-financing and AI-based systems, empowering farmers to unlock the potential of AI. With Jiva, farmers can enhance crop health, streamline operations, and increase profitability, ushering in a promising future for agriculture. AI-powered systems like Jiva offer personalized recommendations for optimal farming practices, resulting in higher crop yields and improved profitability [126].
@@
  • See and Spray: The Blue River Technology ‘See and Spray’ system uses computer vision to identify individual plants and selectively apply herbicides, reducing the amount of chemicals needed and improving crop yields [127].
@@
  • FarmX app: FarmX is an AI-powered app that helps farmers optimize their irrigation and fertilizer usage. The app uses ML algorithms to analyze data from various sources, such as soil moisture sensors, weather data, and satellite imagery. With this information, it generates recommendations for farmers on when and how much water and fertilizer to apply to their crops. The app can also provide alerts if there are any issues with irrigation, such as leaks or blockages. By optimizing irrigation and fertilizer usage, the FarmX app helps farmers improve crop yields while reducing water and fertilizer waste.
AI and robotics are revolutionizing the agricultural industry and helping farmers to address the challenges they face. As technology continues to advance, we can expect even more innovative solutions to emerge in the future.
8 Top 10 AI and Robotics Companies/Start-Up Disrupting Agriculture
@@
The use of AI and robotics in agriculture has been gaining momentum in recent years, and some companies have emerged as global leaders in this field. The top 10 global companies using AI and robotics in agriculture, based on their valuation, are Benson Hill Biosystems, Iron Ox, FarmWise, Carbon Robotics, CropIn, Blue River Technology, Aerobotics, Prospera, AgroScout, and Root AI. The details about this company is provided in Table 8. These companies have developed various technologies, including ML, computer vision, and robotics, to optimize crop growth, monitor plant health, and detect diseases early. By leveraging these technologies, they are helping farmers make data-driven decisions and achieve higher yields while reducing their environmental footprint. Overall, the emergence of these companies is a testament to how AI and robotics can revolutionize agriculture and contribute to a more sustainable and efficient food system.
@@
Table 8: Top 10 global Companies Using AI and Robotics in Agriculture based on their valuation.
Rank Name of Company Headquarters Founded Year Service Website
1 Benson Hill Biosystems USA 2012 Uses ML for analyzing crop gene influence https://bensonhillbio.com/
2 Iron Ox USA 2015 Uses robotics and AI to ensure optimal plant growth https://www.ironox.com
3 FarmWise USA 2016 Builds adaptive robots to farm individual plants https://farmwise.io/
4 Carbon Robotics USA 2018 Makes advanced robotics accessible to improve crop yield https://carbonrobotics.com/
5 CropIn India 2010 Helps with remote sensing, weather advisory, scheduling and monitoring farm activities, educating farmers, and monitoring crop health http://www.cropin.com/
6 Blue River Technology USA 2011 Uses computer vision and robotics to identify lettuce plants with precision http://bluerivertechnology.com
7 Aerobotics South Africa 2014 Uses aerial imagery and ML to help farmers identify pests and diseases early https://www.aerobotics.com/
8 Prospera Israel 2014 Develops computer vision technologies to monitor and analyze plant health and provide actionable insights to growers https://www.prospera.ag/
9 AgroScout Israel 2017 Uses data from multiple sensors, ML, and deep learning to detect and monitor diseases in the field http://agro-scout.com/
10 Root AI USA 2017 Aims to create intelligent robots to help growers build the farms of the future https://root-ai.com/
9 Challenges and Limitations of Machine Learning in Agriculture
@@
While ML and IoT applications [128] in smart farming are changing the way work is done in the agricultural sector, there are still some limitations to the implementation of AI in agriculture that require further research. Few challenges and limitations include as follows:
@@
Additionally, the ownership and control of the data retrieved, stored, and used by the AI system is another issue that needs to be addressed. Some other challenges and limitations of AI in agriculture include: U
@@
  • Lack of technical expertise: The lack of technical expertise among farmers and rural communities can make it difficult to implement and maintain AI technologies in agriculture.
  • Limited awareness: Limited awareness among farmers about new technologies and ethical concerns surrounding AI in agriculture is another barrier to adoption.
@@
10 Future Directions of Machine Learning in Agriculture
@@
  • IoT data: The reliability, consistency, and volume of agricultural data may be compromised due to equipment malfunctions, network node failures, post-processing errors, or pest and disease outbreaks. These data gaps can negatively impact the accuracy of calculations and hinder the effectiveness of IoT applications in agriculture [144].
@@
  • Crop health monitoring: Conventional crop health monitoring techniques are laborious and largely categorical. However, companies are now competing to develop hyperspectral imaging and 3D laser scanning technologies that can enhance precision and accuracy by collecting vast amounts of data. By using deep learning ML techniques, an alert-based system can be developed to improve crop protection [147].
11 Conclusion
The agriculture sector continues to be a crucial priority, supporting the livelihoods of around 50% of the global population. As the potential for the adoption of AI in agriculture becomes increasingly apparent, the gap between current agricultural practices and AI adoption is rapidly closing. In this study, we have conducted an exhaustive review of ML applications in agriculture, providing a comprehensive understanding of the intersection between ML and agriculture.
To establish the foundations, we have explored the basic principles of ML and its relevance to agriculture. In addition, we have presented case studies that exemplify practical applications of AI in the agricultural domain, showcasing how it is used to address real-world challenges and drive significant changes. Furthermore, we have examined the top ten agro-tech companies, highlighting their innovative use of AI and ML to solve pressing issues faced by farmers and drive transformative advancements in the industry. Despite the promising potential, the adoption of AI in agriculture comes with certain challenges and limitations. We have discussed these challenges and limitations in detail, shedding light on the obstacles that need to be overcome for a successful transition. Looking ahead, we have projected the future directions of ML in agriculture through a series of point-wise projections. By considering global demand, infrastructure availability and affordability, as well as national and international initiatives, we have outlined the fundamental changes that are already underway and driving the transformation of the agricultural sector.
Inevitably, the future of agriculture belongs to early adopters of AI. As the gap between current practices and the full potential of AI in agriculture continues to shrink, it becomes evident that embracing AI technologies is essential for sustained progress and success in the industry. By being at the forefront of AI adoption, these early adopters will play a crucial role in shaping the future of agriculture and ensuring its continued growth, sustainability, and prosperity.
Acknowledgments
The authors would like to express their gratitude to the Ministry of Human Resource Development (MHRD) of India for providing financial assistance to the first author through grant number MHR01-23-200-428, and the Council of Scientific & Industrial Research (CSIR) of India for supporting the second author with grant number 09/143(0989)/2019-EMR-I during this research.

### v2

@@ 6.1 Case Studies
The use of AI in agriculture has enormous potential to revolutionize the way farming is conducted, from improving crop management to reducing costs and simplifying tasks for farmers. This potential is being realized by a number of start-ups and applications Here are some examples of how AI is being practically utilized (Table 8) in agriculture:
@@
  • Kisan GPT: It launched on March 15, 2023, is an AI chatbot that caters to the underserved agriculture sector of India. It provides guidance on various topics such as irrigation, pest control, and crop cultivation. Unlike other popular AI chatbots where users are required to type their queries, Kisan GPT offers the option of voice input using a microphone. It supports 10 languages including Hindi, English, Gujarati, and Marathi, as well as foreign languages like Japanese, Portuguese, and Indonesian. With India being an agricultural powerhouse, this AI chatbot could be a game-changer for the agriculture sector. Farmer-friendly and accessible, it bridges the information gap between farmers and experts, aiding decision-making and enhancing crop management. With AI-powered chatbots, like Kisan GPT, farmers can achieve greater efficiency and make informed choices, benefiting India’s agricultural prowess [184].
@@
  • Jiva: Jiva was created by Olam International, a global agribusiness headquartered in Singapore. While it has operations in India too. Jiva provides a range of services, including micro-financing and AI-based systems, empowering farmers to unlock the potential of AI. With Jiva, farmers can enhance crop health, streamline operations, and increase profitability, ushering in a promising future for agriculture. AI powered systems like Jiva offer personalized recommendations for optimal farming practices, resulting in higher crop yields and improved profitability [185].
@@
  • See and Spray: It is a US based system introduced by Blue River Technology. The ’See and Spray’ system uses computer vision to identify individual plants and selectively apply herbicides, reducing the amount of chemicals needed and improving crop yields [186].
@@
  • xFarm: xFarm Technologies is a Swiss-Italian company headquartered in Switzerland, with strong operations in Italy. xFarm is an AI-powered app that helps farmers optimize their irrigation and fertilizer usage. The app uses ML algorithms to analyze data from various sources, such as soil moisture sensors, weather data, and satellite imagery. With this information, it generates recommendations for farmers like when, how much water and fertilizer to apply to their crops. The app can also provide alerts if there are any issues with irrigation, such as leaks or blockages. By optimizing irrigation and fertilizer usage, the xFarm app helps farmers improve crop yields while reducing water and fertilizer waste. [187]
Table 8: Analystical analysis of AI-driven Agricultural Solutions in Case Studies
Case Study Name Data Sources Integrated AI/ML Techniques Used Impact on Agriculture and Farmers
Kisan GPT Weather data, expert recommendations, government advisories, farmer queries $\bullet$ Reinforcement Learning from Human Feedback (RHFL)
$\bullet$ LLMs
$\bullet$ NLP
$\bullet$ Speech Recognition
Real-time multilingual chatbot, removing the barrier of reading and language for farmers, aiding decision-making with expert-backed insights.
Jiva Crop yields, soil health, financial records, farm productivity reports $\bullet$ Predictive Analytics
$\bullet$ Decision Trees
$\bullet$ Financial Risk Modeling
$\bullet$ AI sensors
Provides AI-based farming and micro-financing recommendations, improving financial sustainability. Offers free personalized agronomic advice, cash advances for farm inputs, and facilitates high-quality input purchasing and crop sales.
See and Spray High-resolution camera images, GPS mapping, historical weed infestation patterns, weed data $\bullet$ Deep Learning
$\bullet$ Computer Vision
$\bullet$ Sensor Fusion
Selective herbicide spraying, reducing chemical usage, improving weed control, and increasing crop yield.
xFarm Soil moisture sensors, weather forecasts, satellite imagery $\bullet$ ML-Based Predictive Models
$\bullet$ Geospatial Analytics
$\bullet$ Anomaly Detection
Optimized irrigation and fertilizer management, reducing resource waste and improving yield through data-driven recommendations.
6.2 Top 10 AI and Robotics Companies/Start-Up Disrupting Agriculture
@@
The use of AI and robotics in agriculture has gained momentum in recent years, and some companies have emerged as global leaders in this field. The top 10 global companies using AI and robotics in agriculture, according to their 2023 valuation, are Benson Hill Biosystems, Iron Ox, FarmWise, Carbon Robotics, CropIn, Blue River Technology, Aerobotics, Prospera, AgroScout, and Root AI. The details of this company are provided in Table LABEL:tab:agriculture-companies. These companies have developed various technologies, including ML, computer vision, and robotics, to optimize crop growth, monitor plant health, and detect diseases early. Using these technologies, they help farmers make data-driven decisions and achieve higher yields while reducing their environmental footprint. In general, the emergence of these companies is a testament to how AI and robotics can revolutionize agriculture and contribute to a more sustainable and efficient food system.
@@
Table 9: Top 10 Global Companies Using AI and Robotics in Agriculture Based on Their 2023 Valuation.
Serial No. Company Headquarters Year Founded AI/ML Techniques Used Data Sources Integrated Impact on Agriculture Website
1 Benson Hill Biosystems USA 2012 $\bullet$ Machine Learning (ML) for crop genome analysis Genetic sequencing data, soil properties, crop traits Enhances crop breeding efficiency through AI-driven genetic insights. https://bensonhillbio.com/
2 Iron Ox USA 2015 $\bullet$ AI-powered robotics
$\bullet$ Reinforcement Learning
Environmental sensors, soil moisture, plant health data Reduces water and fertilizer use while optimizing plant growth. https://www.ironox.com
3 FarmWise USA 2016 $\bullet$ Computer Vision
$\bullet$ ML for robotic weeding
Camera images, soil condition data, weed growth patterns Lowers herbicide use by automating precision weeding. https://farmwise.io/
4 Carbon Robotics USA 2018 $\bullet$ Deep Learning
$\bullet$ AI-driven autonomous weeding robots
Multi-spectral imaging, GPS farm mapping Eliminates weeds without chemicals, improving sustainability. https://carbonrobotics.com/
5 Blue River Technology USA 2011 $\bullet$ Computer Vision
$\bullet$ AI-based precision agriculture
High-resolution imaging, real-time plant analysis Reduces chemical use with plant-specific treatments. http://bluerivertechnology.com
6 Root AI USA 2017 $\bullet$ AI-powered autonomous robots for harvesting Sensor-based crop monitoring, fruit ripeness detection Improves farm efficiency using robotic automation. https://root-ai.com/
7 CropIn India 2010 $\bullet$ AI-driven satellite imaging
$\bullet$ Predictive Analytics
Remote sensing, weather forecasts, soil health data Supports precision farming through risk management and remote crop monitoring. http://www.cropin.com/
8 Aerobotics South Africa 2014 $\bullet$ Deep Learning for aerial imaging Drone-captured images, temperature data, vegetation indices Detects crop diseases early, minimizing yield losses. https://www.aerobotics.com/
9 Prospera Israel 2014 $\bullet$ ML
$\bullet$ Computer Vision for crop monitoring
Farm camera networks, soil health analytics Enhances yield with real-time crop health analysis. https://www.prospera.ag/
10 AgroScout Israel 2017 $\bullet$ Multi-Sensor Data Fusion
$\bullet$ Deep Learning
Satellite imagery, IoT sensor data, weather patterns Enables early-stage pest and disease detection. http://agro-scout.com/
7 Publicly available dataset related to Agriculture
Training an ML model for experiments is highly dependent on the quality and structure of the dataset employed. Frequently, these datasets comprise raw data that must be preprocessed before model training. In addressing RQ-5, this section highlights several publicly available datasets to assist the research community in selecting appropriate data for their studies. These datasets are classified into three categories—weed control, fruit detection, and miscellaneous—as detailed in Table 10, Table 11, and Table 12, respectively.
The available information for these datasets is summarized in terms of Modality and Annotation. Modality refers to the different types of data available, such as multispectral, RGB, etc. Annotation refers to the task of labeling images. Specifically, the pixel-level annotation [211] also known as semantic segmentation, is the leveling of each pixle of an image with a specific class. This annotation provides a detailed understanding of the image by identifying the exact boundaries and shapes of the objects. On the other hand, image level annotation [212], which is defined as labeling the entire image with a single class or a set of classes. This type of annotation gives an understanding of an image’s content without the exact location of objects.
Table 10: Public Datasets Dedicated to Weed Control.
No. Dataset Name Modality Annotation Ref.
1 CWFI dataset Multispectral Pixel level [239]
2 Carrot-Weed RGB Pixel level [240]
3 Plant seedlings RGB Image level [241]
4 Grass-Broadleaf RGB Patch level [242]
5 Sugar Beets 2016 Multimodal Image level [243]
6 Synthetic SugarBeet Weeds RGB Pixel level [244]
7 WeedNet Multispectral Image level [245]
8 Joint stem detection Multispectral+RGB Pixel level [246]
9 Leaf counting RGB Image level [247]
10 Weed Map Multispectral Pixel level [248]
11 DeepWeeds RGB Image level [249]
12 Crop weed discrimination Multispectral Pixel level [250]
13 Early crop weed RGB Image level [251]
14 Ladybird Cobbitty Brassica Multimodal No annotations [252]
15 Open Plant Phenotype Database (OPPD) RGB Image level, bounding box [253]
Table 11: Public Datasets Dedicated to Fruit Detection.
No. Dataset Modality Annotation Ref.
1 DeepFruits RGB Bounding box [254]
2 Orchard Fruit RGB Bounding box, circle [255]
3 Date Fruit RGB Image level [256]
4 KFuji RGB-DS RGB-D Bounding box [257]
5 MangoNet RGB Pixel level [258]
6 MangoYOLO RGB Bounding box [259]
7 WSU apple dataset RGB Bounding box [260]
8 Fuji-SfM RGB Bounding box [261]
9 LFuji-air dataset LiDAR Bounding box [262]
10 MinneApple RGB Pixel level [263]
Table 12: Miscellaneous Public Datasets.
No. Dataset Modality Annotation Ref.
1 3D Broccoli RGBD No annotation [264]
2 Apple Trees RGBD No annotation [265]
3 Capsicum Annuum RGB Pixel level [266]
4 Fruit flower dataset RGB Pixel level [267]
5 Sugarcane billets RGB Image level [268]
6 Maize disease RGB Line level [269]
7 DeepSeedling RGB Bounding box [270]
8 GrassClover RGB Pixel level [271]
9 Oil radish growth RGB Pixel level [272]
10 University of Arkansas, Plants Dataset Info not available Info not available [275]
11 EPFL, Plant Village Dataset Info not available Info not available [276]
12 Leafsnap Dataset Info not available Info not available [277]
13 LifeCLEF Dataset Info not available Info not available [278]
14 PASCAL Visual Object Classes Dataset Info not available Info not available [279]
15 Africa Soil Information Service (AFSIS) dataset Info not available Info not available [280]
16 UC Merced Land Use Dataset Info not available Info not available [281]
17 MalayaKew Dataset Info not available Info not available [282]
18 Crop/Weed Field Image Dataset Info not available Info not available [283]
19 University of Bonn Info not available Info not available [284]
20 Flavia Leaf Dataset Info not available Info not available [285]
21 Syngenta Crop Challenge 2017 Info not available Info not available [286]
8 Challenges and Future Research Directions
8.1 Challenges and Limitations of Machine Learning in Agriculture
@@
Although ML and IoT applications [209] in smart agriculture are changing the way work is done in the agricultural sector, there are still some limitations to the implementation of AI in agriculture that require further research. Few challenges and limitations include as follows:
@@
Additionally, the ownership and control of the data retrieved, stored, and used by the AI system [216] is another issue that needs to be addressed. Some other challenges and limitations of AI in agriculture include:
@@
  • Bridging the Awareness Gap: Limited awareness among farmers about new technologies and ethical concerns surrounding AI in agriculture is another barrier to adoption.
@@
8.2 Challenges in Data Fusion for Agricultural Applications/Heterogeneous Data Fusion Challenges
As mentioned in the dataset section, various types of data are used in the agriculture field, originating from multiple sources such as RGB images from cameras, remote sensing, UAV images, satellite imagery, IoT sensors, weather models, soil quality reports, and economic datasets. The term data fusion refers to the process of combining data from multiple sources to produce more accurate, consistent, and concise information than that provided by individual sources [17]. While machine learning (ML) applications have advanced precision agriculture, data fusion challenges remain a significant barrier to effective decision-making.
  • Data Format Incompatibility: Agricultural data is available in multiple formats, making data fusion complex. For instance, satellite, drone, and camera images are stored as raster formats (JPEG, PNG, etc.), whereas meteorological data is represented as time-series or numerical data (JSON, SQL, CSV, etc.). Combining these diverse formats requires specialized techniques to align their structure and ensure compatibility for analysis [9, 11].
  • Spatial and Temporal Misalignment: Different data sources operate at different time scales and resolutions. For instance, data from IoT sensors is collected in real time, while satellite images are taken approximately every 5 to 10 days. Therefore, fusing these types of data is impractical without preprocessing and aligning them to the same time scale [10]
  • Computational Complexity: Working with deep learning models on high-dimensional and large-scale datasets often requires GPUs. For instance, fusing time-series data with spatial pixel-based data for agricultural applications can be achieved using a hybrid architecture that combines CNN and LSTM models, in this scenario, data fusion further increases computational complexity [12].
  • Network Latency and Connectivity Issues: Smart farming practices require real-time decision-making, which heavily depends on IoT-based agricultural systems relying on wireless connections (Wi-Fi or strong internet connectivity). However, poor connectivity in rural areas leads to delayed data transmission, which, in turn, affects the real-time processing and predictions of ML models [18].
  • Missing and Conflicting Data from Multiple Sources: One of the major challenges in applying ML models to smart farming is handling missing, inconsistencies, and conflicting data, all of which negatively effect real time agricultural decision-making. Various factors contribute to these issues, including weather conditions, misplacement of IoT sensors, and other environmental influences. For example, satellite data may be unavailable due to cloud cover, resulting in missing data. Additionally, different data sources may provide inconsistent measurements—for instance, soil pH readings from IoT sensors might differ from lab-based data—making it difficult for ML models to accurately assess real conditions [19].
8.3 Future Directions of Machine Learning in Agriculture
@@
  • IoT and Multi sensors data : The reliability, consistency, and volume of agricultural data may be compromised due to equipment malfunctions, network node failures, post-processing errors, or pest, disease outbreaks and data from multi sensors . These data gaps can negatively impact the accuracy of calculations and hinder the effectiveness of IoT sensors applications in agriculture [118].
@@
  • Crop health monitoring: Conventional crop health monitoring techniques are laborious and largely categorical. However, companies are now competing to develop hyperspectral imaging and 3D laser scanning technologies that can enhance precision and accuracy by collecting vast amounts of data. By using deep learning [214, 215] techniques, an alert-based system can be developed to improve crop protection [116].
  • Model Interpretation: ML models are often considered black boxes, making it challenging for researchers and practitioners to understand why and how they make decisions [13]. These challenges raise concerns about the trustworthiness of their predictions in agricultural applications for farmers. Developing interpretable models for agricultural applications can help mitigate these issues [14].
  • Incorporating Generative Adversarial Networks: Generative Adversarial Networks (GANs) can address data scarcity and imbalance in agricultural datasets for ML and deep learning models. Agricultural datasets often lack diversity in environmental conditions, leading to overfitting or underfitting in ML and DL models [15]. Beyond image generation, GANs offer solutions for multi-sensor and multi-source data fusion. Future research can explore domain-specific GAN architectures to enhance their applicability in agriculture [16].
9 Discussion and Conclusion
This review provides a comprehensive analysis of over 70 research articles on ML applications in agriculture, addressing the research objectives outlined at the beginning of the study. In doing so, it highlights key advancements and real-world implementations, such as Kisan GPT, Jiva, CropIn, and Blue River Technology, while offering a foundational resource for new researchers.
Addressing RO-1, the review investigates recent advancements and applications of ML across all agricultural stages—from pre-harvesting (e.g., soil classification, weather prediction, seed testing, and disease detection) to harvesting (e.g., fruit detection and robotic picking) and post-harvesting processes (e.g., quality assessment and food security). Tables LABEL:tab:long and 3 succinctly summarize these contributions, providing a quick reference to key developments in each domain.
In line with RO-2, the study demonstrates how ML is integrated with agricultural data and data fusion techniques to extract valuable insights. It discusses various methodologies for data fusion, including ensemble learning, Bayesian inference, hybrid models, and deep learning-based feature fusion, that collectively enhance decision-making and productivity in precision agriculture.
Addressing RO-3, a detailed bibliometric analysis is presented, revealing a substantial growth in research output over recent years. Figures 7, 8, and 9 illustrate research trends, global contributions, and institutional collaborations. The statistical findings, supported by an exponential trendline (with an $R^{2}$ value of 0.993) and publication distribution data (Table 7), underscore the dynamic and vibrant nature of ML research in agriculture.
In response to RO-4, real-world applications and case studies are reviewed to bridge the gap between academic research and practical implementation. The discussion of case studies and the analysis of approximately ten leading AI and robotics companies in agriculture (detailed in Table 8) showcase how these technologies are revolutionizing farming practices on a global scale.
Finally, addressing RO-5, the review lists several publicly available datasets categorized into weed control, fruit detection, and miscellaneous datasets. These datasets, detailed in Tables 10, 11, and 12, serve as a valuable resource for practitioners aiming to build ML models for agricultural applications.
Collectively, these findings not only highlight current research trends and technological advancements but also lay a robust foundation for future collaborative efforts and innovations in the application of ML in agriculture. While this review does not delve into the technical intricacies of individual ML algorithms or the broader socio-economic impacts of agricultural AI, it provides comprehensive insights that will guide future research directions and practical implementations in the field.
Acknowledgement
The author expresses her gratitude to the Ministry of Human Resource Development (MHRD), Government of India, for providing financial assistance for her research and ANID FONDECYT project 123112, ANID PIA/PUENTE AFB230002, Chile.